OpenStudy (anonymous):

Find the equation of the line that passes through the points (-2, 1) and (1, -2).

OpenStudy (mathmale):

I'd suggest a 2-step process: (1) find the slope of the line through these two points. Call it m. (2) choose one of the given points. Substitute its coordinates and m into the point-slope form of the equation of a straight line. (3) simplify the result.

OpenStudy (texaschic101):

find the slope first by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (-2,1) x1 = -2 and y1 = 1 (1,-2) x2 = 1 and y2 = -2 now we sub slope(m) = (-2 -1) / (1 - (-2) slope(m) = -3/3 which equals -1 now we will use y = mx + b slope(m) = -1 (-2,1) x = -2 and y = 1 now sub 1 = -1(-2) + b 1 = 2 + b 1 - 2 = b -1 = b so your equation is : y = -x - 1

OpenStudy (mathmale):

KM: You could check this result by substituting the x- and y-coordinates of either (or both) points and deciding whether the resulting equation is true or false.

OpenStudy (texaschic101):

or, like mathmake said, you could have used point slope formula instead of slope intercept formula

OpenStudy (texaschic101):

oops...mathmale...sorry bout that